b) f(x) = −3x2 + 7

c) f(x) = (x + 1)∕(x + 2)


1
Expert's answer
2021-11-17T07:03:49-0500

Question is incomplete.

Let us ask for "f(0)" :

Solution:

(b)

"f(x) = −3x^2 + 7\\
f(0)=-3(0)^2+7=0+7=7"


(c)

"f(x) = (x + 1)∕(x + 2)\\
f(0)=(0+1)/(0+2)=1/2"